OpenStudy (anonymous):

find the value of c if sigma n = 2 to infinity of (1+c)^-n = 2

OpenStudy (sirm3d):

simplify first with \[y=\frac{1}{1+c}=(1+c)^{-1}\] \[\sum_{n=2}^\infty (1+c)^{-n}=2\\\sum_{n=2}^\infty y^n=2\] \[\large \frac{1}{1-y}=\sum_{n=0}^\infty y^n=1+y+(y^2+y^3+\cdots)=1+y+\sum_{n=2}^\infty y^n\]

OpenStudy (sirm3d):

\[\large \sum_{n=2}^\infty y^n=\frac{1}{1-y}-1-y\] \[\large {2=\sum_{n=2}^\infty y^n=\frac{1}{1-y}-1-y\\2=\frac{1}{1-y}-1-y}\] solve y, then solve c